OpenStudy (birdy258):

what is the vaule of f^1 (f(x))

OpenStudy (anonymous):

????????

OpenStudy (birdy258):

that's what the question was

OpenStudy (anonymous):

does it say what x is

OpenStudy (birdy258):

no it wants the vaule of what i put in its 0,1 ,x,x^-1

OpenStudy (anonymous):

oh dang man i dont know what to say about that im sorry what about you @phi

OpenStudy (birdy258):

man i need help

OpenStudy (anonymous):

@welshfella can you help out on this one

OpenStudy (birdy258):

nvm it was x i guessed and failed the test

OpenStudy (welshfella):

i can give you a direct answer but that s not allowed basiccally what we have here is the x of a function being replaced by the inverse of the function.

OpenStudy (welshfella):

yes it is x

OpenStudy (anonymous):

thanx @welshfella

OpenStudy (welshfella):

lets try a simple one to illustrate f(x) = 2x ] y = 2x x = y / 2 so f-1(x) = x/2 f-1(fx) = 2 * x/2 = x

OpenStudy (anonymous):

you lost me i am good at math but that is confusing

OpenStudy (phi):

inverse functions can be confusing. but if you have a pair (x,y) we say we "put x into" some function, and it "returns" y (maybe like a vending machine: money in, candy out" the "inverse function" does it the other way round: you put in y and get out x so if you do \[ f^{-1}(f(x)) \] the f(x) "returns" y \[ f^{-1}(y) \] and if you put y into the inverse function, you get back x \[ x= f^{-1}(y) \] in other words \[ x= f^{-1}(f(x)) \] It gets confusing, because people always use "x" as the input variable even when writing down the inverse function: \( f^{-1}(x) \) so you have to remember that "x" is the "y" of the f(x) function. *very confusing*